OpenStudy (anonymous):

3x + y = 6 x - y = 6

OpenStudy (anonymous):

Add the equations together: 3x + x + y - y = 6 + 6 Simplified: 4x = 12 Simplified: x = 3 Now we can plug in x into one of the first equations to find y 3(3) + y = 6 Simplify: 9 + y = 6 Simplify: y = -3